FREE SUBPRODUCTS AND FREE SCALED PRODUCTS OF II_{1}-FACTORS

KEN DYKEMA

Abstract. The constructions of free subproducts of von Neumann algebras and free scaled products are introduced, and results about them are proved, including rescaling results and results about free trade in free scaled products.